# VIPER: Process-aware Evaluation for Generative Video Reasoning

### v1

@@ Abstract
Recent breakthroughs in video generation have demonstrated an emerging capability termed Chain-of-Frames (CoF) reasoning, where models resolve complex tasks through the generation of continuous frames. While these models show promise for Generative Video Reasoning (GVR), existing evaluation frameworks often rely on single-frame assessments, which can lead to outcome-hacking, where a model reaches a correct conclusion through an erroneous process. To address this, we propose a process-aware evaluation paradigm. We introduce VIPER, a comprehensive benchmark spanning 16 tasks across temporal, structural, symbolic, spatial, physics, and planning reasoning. Furthermore, we propose Process-outcome Consistency (POC@r@r), a new metric that utilizes VLM-as-Judge with a hierarchical rubric to evaluate both the validity of the intermediate steps and the final result. Our experiments reveal that state-of-the-art video models achieve only about  20% POC@1.0 and exhibit a significant outcome-hacking. We further explore the impact of test-time scaling and sampling robustness, highlighting a substantial gap between current video generation and true generalized visual reasoning. Our benchmark will be publicly released.
VIPER: Process-aware Evaluation for Generative Video Reasoning
@@ 1 Introduction
Recent breakthroughs in video generation have enabled models to produce highly realistic, consistent, and extended sequences of real-world scenes OpenAI (2025b); Google (2025); Chen et al. (2025b). While these advancements demonstrate immense potential for artistic creation and digital media, recent research suggests that their impact extends far beyond content generation: video models are beginning to exhibit generalized visual reasoning capabilities. Specifically, by generating continuous sequences of frames, video models can depict the step-by-step resolution of tasks they were not explicitly trained for, ranging from fundamental visual processing (e.g., super-resolution) to high-level logical reasoning (e.g., maze solving). This emerging capability is termed “Chain-of-Frames” (CoF) reasoning Wiedemer et al. (2025).
Figure 1: POC@1.0 performance overview of representative video model on VIPER across 6 domains.
@@
To better understand this phenomenon, researchers have introduced various Generative Video Reasoning (GVR) tasks. These tasks typically require a model to generate a video that completes a specific reasoning goal based on an initial image. While current evaluations show that video models possess promising reasoning potential, we identify several critical flaws in existing GVR assessment frameworks. A primary issue is that existing evaluation data and strategies provide insufficient focus on the procedural nature of video. Since a video consists of continuous frames, it possesses an inherent temporal and procedural attribute. However, existing evaluation strategies typically sample the last or the best-matched frame for assessment. This approach evaluates the correctness of the outcome while disregarding the validity of the reasoning process itself. According to our empirical study in Section 2, current video models produce Outcome-hacking, where model generates videos that reach a correct conclusion through an erroneous process, Existing evaluation metrics based on single frame often misclassify these cases as correct, leading to an overestimation of the models’ true reasoning capabilities. Moreover, in terms of evaluation metrics, most current work either repurposes metrics from other domains or relies on generic VLM scoring, lacking a design specifically tailored for GVR tasks.
@@
To address these issues, we propose a process-aware evaluation paradigm for GVR. First, we introduce VIPER (VIdeo Process Evaluation for Reasoning tasks), a comprehensive benchmark consisting of 16 tasks across 6 domains, namely temporal, structural, symbolic, spatial, physics, and planning reasonging, totally 309 items. To ensure a reliable evaluation of generated content, we propose a new metric, Process-outcome Consistency (POC@r@r). POC@r@r takes frames sampled from the entire video at a sampling rate $r$ as input. For any given video, the POC score is determined by both its process consistency (PC) and outcome consistency (OC). OC requires at least one frame to fulfill the explicit task goal, while PC requires all sampled frames to adhere to the task’s process-level constraints. A video is considered correct if and only if it fulfills both OC and PC. In practice, we follow the VLM-as-Judge paradigm to conduct these evaluations, utilizing a hierarchical rubric designed to facilitate accurate and extensible assessment.
@@
Our experimental results demonstrate that even state-of-the-art video models only achieve approximately 30% POC@1.0, while exhibiting an outcome-hacking rate of around 20%. these findings indicate that a significant gap remains before current video models can be considered truly generalizable visual reasoners, particularly regarding the consistency between their processes and outcomes. We also investigate the effects of test-time scaling, a prevailing technique in language reasoning, on video models and find it yields notable performance improvements. Furthermore, we study the robustness of the POC under different sampling rate r. Finally, we validate the robustness of our evaluation method through human-model consistency checks and summarize several common reasoning failure modes in existing models.
@@ 2.1 Experiment Design
Figure 3: Overview of VIPER. VIPER consists of 16 tasks from 6 domains focusing on different reasoning abilites.
@@ 3.1 Generative Video Reasoning
    Implicit constraints ($c$): Latent rules not directly specified in the prompt but inherent to the task (e.g., “do not cross the walls of the maze”).
@@ 3.2 Evaluation Dimension
To comprehensively evaluate the reasoning capabilities of video generation models, we propose six evaluation dimensions that focusing different aspect. Each dimension contains multiple tasks, as illustrated in Figure 3. The detailed introduction of each task is presented in the Appendix A.
@@
This domain focuses on evaluating the model’s ability to accurately track and represent the dynamic evolution of objects or scenes over time. We design the following tasks: (1) Object Moving: simulate the movement of simple geometric shapes to the target position. (2) World Memory: simulate zooming in and out of an image while keeping the content unchanged.
@@
This domain focuses on highly structured tasks that require the model to follow strict, predefined rules and constraints. We include the following reasoning tasks: (1) Maze: navigate from the start to the goal without violating maze constraints. (2) Checkmate-in-One: find a single legal move that delivers checkmate under standard chess rules. (3) Tic-Tac-Toe: choose a legal move to win or block the opponent under the game rules. (4) Sudoku: fill the grid so that all rows, columns, and subgrids have no duplicated digits.
@@
This domain emphasizes the model’s ability to reason using symbolic information (e.g., text tokens) within the generated video. We collect a variety of tasks: (1) Math Reasoning: illustrate the process of solving a math problem. (2) Knowledge Reasoning: solve problems based on world knowledge. (3) Multimodal Reasoning: solve problems involving both text and image information.
@@
This domain evaluates the model’s ability to understand and reason about spatial relationships, geometric transformations, and the relative positioning of objects. We collect the following tasks: (1) Dice Rolling: simulate the rolling of a die along a specified trajectory. (2) Cube Rotation: simulate the rotation of a composite structure composed of multiple cubes. (3) Image Puzzle: reassemble a set of shuffled sub-images into their original configuration.
@@ 3.3 Data Collection
Each sample in our dataset is formulated as a quadruple $\{i,p,c,g\}$, where $i$ represents the input image, $p$ denotes the task prompt, $c$ encapsulates the process constraints, and $r$ serves as the ground-truth reference (text, images, or videos).
@@
Input images are curated from diverse sources, including web collection, programmatic generation via Python, synthesis using image generation models, and selection from existing datasets.
@@
To further facilitate accurate evaluation, we provide references that depict the correct process for each sample. Depending on the task, references may vary: for code-rendered tasks, the reference is the ground-truth rendered video; for scenarios where intermediate process data is unavailable, we provide the image of the solved state; for symbolic reasoning tasks, the reference consists of both the textual reasoning steps and the final answer.
@@ 4.1 Process-outcome Consistency
Formally, let $V=\mathcal{M}(i,t)$ denote the video generated by model $\mathcal{M}$. Instead of select the last or best frame for evaluation, POC uniformly sample multiple frames from $V$ at a frame rate of $r$ (denoted as $\text{POC}@r$), obtaining the frame set $\hat{V}_{r}=\{f_{k}\}_{k=1}^{N}$, where $N$ is the total number of sampled frames. To determine whether a frame adheres to the text condition, we employ a VLM as the judge model $\mathcal{J}$. Specifically, POC@r@r is formulated as:
@@
Table 2: Performance of mainstream video models VIPER arcoss 16 tasks within . We report the POC@1.0 for each domain and summarize the average OC, POC and hacking rate. Best performance is bolded.
@@ 4.2 Hierarchical Rubric
To facilitate accurate consistency judgment, we design a hierarchical rubric for $\mathcal{J}$. Specifically, the rubric is structured into three distinct levels:
@@
This level defines the task context, evaluation criteria, and methodology. Adopting the popular test-time scaling paradigm Guo et al. (2025a), the system prompt requires the model to first generate the analysis of process and outcome consistency within <think> tags, and subsequently provide the decision for PC, OC and POC within <answer> tags.
@@
This level provides a concise overview of the domain to which the evaluated task belongs, highlighting the specific evaluation focus within that field. Notably, all tasks within the same domain share a unified domain introduction.
@@
This level aims to provide fine-grained, task-specific evaluation criteria. These criteria are reorganized from the previously collected process constraints $c$ and the reference $g$.
Collectively, this hierarchical rubric covers from macro-level task definitions to micro-level requirements, providing comprehensive evaluation guidance. Moreover, this design is highly scalable. To support new tasks or domains, we only need to add the corresponding constraints or introductions.
@@ 5.2 Experiment Result
We present the main experiment results in Table 2, from which we summarize following observations. Firstly, all models struggle with VIPER, with POC@1.0 scores consistently below 30. Sora 2 performs the best, achieving a score of 23.3, followed by Veo 3.1 and Wan 2.6, scoring 20.3 and 18.6. The remaining models fail to surpass a score of 10. This indicates that current video generation models are not yet capable of functioning as general visual reasoners. Furthermore, severe outcome-hacking is observed across all models. Notably, Veo 3.1 experiences the most severe hacking, with a rate of 35.8. This suggests that, while these models possess the necessary knowledge for fulfill reasoning goals, they struggle to faithfully represent the reasoning process through video frames. As such, improving process-level reasoning is a crucial step toward advancing model performance. Lastly, the models demonstrate varying strengths across different tasks. Symbolic reasoning, in particular, proves to be a major challenge for most models, which struggle to generate accurate and legible text. However, Sora 2 excels in this domain, achieving a POC of 58.3, which highlights its robust text rendering and reasoning capabilities. In contrast, for tasks like planning and physics which are more closely tied to real-world scenarios, Veo 3.1 outperforms Sora 2, underscoring its superior capabilities in real-world physics simulation.
Takeaway 1 There remains a significant gap between existing video models and universal visual reasoners, especially regarding domains like spatial and physical reasoning and the consistency between the process and outcome.
@@
Table 3: Effect of test-time scaling. We calculate Pass@@1, Pass@@4, Pass@@8 for Veo 3.1 and Sora2 with POC@@1.0.
@@ 5.3 Further Analysis
We investigate whether test-time scaling techniques, which have proven effective in language reasoning, can similarly benefit video reasoning tasks. Pass@$k$ is a widely used metric for evaluating the reasoning capabilities of large models, with recent studies indicating that performance can be significantly enhanced by increasing the number of samples $k$ Chen et al. (2021); Wang et al. (2023). To validate this in the GVR tasks, we sample multiple videos on VIPER using Sora 2 and Veo 3.1 and reporting Pass@1, 4, and 8 with POC@@1.0 in Table 3. As illustrated, scaling the sampling times notably improves performance. For instance, the overall accuracy of Veo 3.1 increases from 20.3% to 43.2%, particularly in challenging domains such as symbolic reasoning. However, this upward trend begins to plateau as $k$ increases, suggesting that test-time scaling alone is insufficient to overcome fundamental reasoning limitations.
@@
Takeaway 2 Scaling sampling times can effectively improve model performance, but it is insufficient to fundamentally enhance the reasoning capabilities of video models.
@@
Figure 4: Human-model alignment under of combinations of raw input and hierarchical rubric.
@@
To further validate the robustness of our evaluation paradigm, we compare the evaluation results from the judge model with those from human experts. As illustrated in Figure 4, by providing the judge model with our hierarchical rubric, its alignment with human evaluators generally exceeds 90%, whereas the raw evaluation achieves only about 60%. These results demonstrate the effectiveness and necessity of our rubric.
Figure 5: Representative failure patterns selected from VIPER.
@@ 5.4 Failure Patterns
By observing the generation results on the VIPER benchmark, we identify several representative failure modes prevalent in current video models. We categorize these patterns below and provide cases in Figure 5, aiming to provide insights and potential directions for future optimization of video models.
@@
We observe that models often struggle to identify implicit constraints within tasks, such as not violating the maze structure when solving the maze or adhering to the movement rules in a Checkmate-in-one scenario. To address this, we investigate whether explicitly providing these constraints as input can improve the model’s performance. We test Veo 3.1 by providing the model with the process constraints $c$. As shown in Table 5, providing these additional information significantly enhances the model’s performance, particularly in domains where constraints are crucial, such as temporal and physics reasoning.
@@
Current video generation models generally produce videos of fixed lengths. However, some tasks, such as Checkmate-in-one, involve reasoning steps that are relatively short and may correspond to only 1-2 seconds of video. Ideally, the model should maintain a still frame after completing the task. However, current models often continue to make unnecessary changes to the content of the frame after the task is completed, disrupting the integrity of the reasoning process.
@@
Unlike general video generation, GVR tasks are more structured and typically require the model to perform highly controlled editing based on the given input. However, current models struggle with defining clear boundaries for editing and often make unnecessary modifications to the image content. These include altering the background, introducing new objects unrelated to the task, or violating the original task conditions.
@@
The readability of text within the generated video is a crucial factor in determining the reasoning quality, particularly in symbolic reasoning tasks where the model is expected to render the solution process in textual form. Currently, most models struggle to generate fluent and readable text passages. As a result, even if the reasoning outcome is correct, the video fails to provide a coherent and interpretable reasoning process.

### v2

# Beyond the Last Frame: Process-aware Evaluation for Generative Video Reasoning

@@ Abstract
Recent breakthroughs in video generation have demonstrated an emerging capability termed Chain-of-Frames (CoF) reasoning, where models resolve complex tasks through the generation of continuous frames. While these models show promise for Generative Video Reasoning (GVR), existing evaluation frameworks often rely on single-frame assessments, which can lead to outcome-hacking, where a model reaches a correct conclusion through an erroneous process. To address this, we propose a process-aware evaluation paradigm. We introduce VIPER, a comprehensive benchmark spanning 16 tasks across temporal, structural, symbolic, spatial, physics, and planning reasoning. Furthermore, we propose Process-outcome Consistency (POC@r@r), a new metric that utilizes VLM-as-Judge with a hierarchical rubric to evaluate both the validity of the intermediate steps and the final result. Our experiments reveal that state-of-the-art video models achieve POC@@1.0 only about  20% and exhibit a significant outcome-hacking. We further explore the impact of test-time scaling and sampling robustness, highlighting a substantial gap between current video generation and true generalized visual reasoning. Our benchmark are released at https://github.com/RUCAIBox/VIPER.
Beyond the Last Frame: Process-aware Evaluation for
Generative Video Reasoning
@@ 1 Introduction
Recent breakthroughs in video generation have enabled models to produce highly realistic, consistent, and extended sequences of real-world scenes OpenAI (2025b); Google (2025); Chen et al. (2025b). While these advancements demonstrate immense potential for artistic creation and digital media, recent research Rotstein et al. (2025); Wiedemer et al. (2025) suggests that their impact extends far beyond content generation: video models are beginning to exhibit generalized visual reasoning capabilities. Specifically, by generating continuous sequences of frames, video models can depict the step-by-step resolution of tasks ranging from fundamental visual processing (e.g., super-resolution) to high-level logical reasoning (e.g., maze solving). This emerging capability is termed “Chain-of-Frames” (CoF) reasoning.
Figure 1: POC@1.0 performance overview of representative video models on VIPER across 6 domains.
@@
To better understand this phenomenon, researchers have introduced various Generative Video Reasoning (GVR) tasks. These tasks typically require a model to generate a video that completes a specific reasoning goal based on an initial image. While current evaluations show that video models possess promising reasoning potential, we identify several critical flaws in existing GVR assessment frameworks. A primary issue is that existing evaluation data and strategies provide insufficient focus on the procedural nature of video. Since a video consists of continuous frames, it possesses an inherent temporal and procedural attribute. However, existing evaluation strategies typically sample the last or the best-matched frame for assessment. This approach evaluates the correctness of the outcome while disregarding the validity of the reasoning process itself. According to our empirical study in Section 2, current video models produce outcome-hacking, where a model generates videos that reach a correct conclusion through an erroneous process. Existing evaluation metrics based on single frame often misclassify these cases as correct, leading to an overestimation of the models’ true reasoning capabilities.
@@
To address these issues, we propose a process-aware evaluation paradigm for GVR. First, we introduce VIPER (VIdeo Process Evaluation for Reasoning tasks), a comprehensive benchmark consisting of 16 tasks across 6 domains, namely temporal, structural, symbolic, spatial, physics, and planning reasoning, comprising 309 items. To ensure a reliable evaluation of generated content, we propose a new metric, Process-outcome Consistency (POC@r@r). POC@r@r takes frames sampled from the entire video at a sampling rate $r$ as input. For any given video, the POC score is determined by both its process consistency (PC) and outcome consistency (OC). OC requires at least one frame to fulfill the explicit task goal, while PC requires all sampled frames to adhere to the task’s process-level constraints. A video is considered correct if and only if it fulfills both OC and PC. In practice, we follow the VLM-as-Judge paradigm to conduct these evaluations, utilizing a hierarchical rubric designed to facilitate accurate and extensible assessment.
@@
Our experimental results demonstrate that even state-of-the-art video models only achieve approximately 20% POC@1.0, while exhibiting an outcome-hacking rate of around 30%. These findings indicate that a significant gap remains before current video models can be considered truly generalizable visual reasoners, particularly regarding the consistency between their processes and outcomes. We also investigate the effects of test-time scaling, a prevailing technique in language reasoning, on video models and find it yields notable performance improvements. Furthermore, we investigate the impact of the sampling rate $r$ on evaluation outcomes. Finally, we verify the reliability of our evaluation method through human-model alignment checks and summarize several common reasoning failure modes in existing models.
@@ 2.1 Experiment Design
Figure 3: Overview of VIPER. VIPER consists of 16 tasks from 6 domains focusing on different reasoning abilities.
@@ 3.1 Generative Video Reasoning
    Implicit constraints ($C$): Rules not directly specified in the prompt but inherent to the task (e.g., “do not cross the walls of the maze”).
@@ 3.2 Evaluation Dimension
To comprehensively evaluate the reasoning capabilities of video generation models, we propose six evaluation dimensions that focusing different aspect. Each dimension contains multiple tasks, as illustrated in Figure 3. The detailed introduction of each task is presented in the Appendix B.
@@
Tasks from this domain focus on evaluating the model’s ability to accurately track and represent the dynamic evolution of objects over time. We design the following tasks: (1) Object Moving: simulate the movement of simple geometric shapes to the target position. (2) World Memory: simulate zooming in and out of an image while keeping the content unchanged.
@@
Tasks from this domain focus on highly structured scenarios that require the model to follow strict, predefined rules and constraints. We include the following reasoning tasks: (1) Maze: navigate from the start to the goal without violating maze constraints. (2) Checkmate-in-One: find a single legal move that delivers checkmate under standard chess rules. (3) Tic-Tac-Toe: choose a legal move to win or block the opponent under the game rules. (4) Sudoku: fill the grid so that all rows, columns, and subgrids have no duplicated digits.
@@
Tasks from this domain emphasize the model’s ability to reason using symbolic information (e.g., text tokens) within the generated video. We collect a variety of tasks: (1) Math Reasoning: illustrate the process of solving a math problem. (2) Knowledge Reasoning: solve problems based on world knowledge. (3) Multimodal Reasoning: solve problems involving both text and image information.
@@
Tasks from this domain evaluate the model’s ability to understand and reason about spatial relationships, geometric transformations, and the relative positioning of objects. We collect the following tasks: (1) Dice Rolling: simulate the rolling of a die along a specified trajectory. (2) Cube Rotation: simulate the rotation of a composite structure composed of multiple cubes. (3) Image Puzzle: reassemble a set of shuffled sub-images into their original configuration.
@@ 3.3 Data Collection
Each sample in our dataset is formulated as a quadruple $\{i,p,C,g\}$, where $i$ represents the input image, $p$ denotes the task prompt, $C$ encapsulates the process constraints, and $g$ serves as the ground-truth reference (text, images, or videos). In total, we collect 309 items, detailed statistics are presented in Appendix A.
@@
Input images are curated from diverse sources, including web collection, Python programs, advanced image generation models, and selection from existing datasets.
@@
To facilitate precise evaluation, we provide a reference $g$ representing the ground-truth process or outcome. The format of $g$ varies by task: it is sampled from the ground-truth video if available; otherwise, it is an image of the solved state. For symbolic reasoning tasks, $g$ comprises the textual reasoning steps and the final answer.
@@ 4.1 Process-outcome Consistency
Formally, let $V=\mathcal{M}(i,t)$ denote the video generated by model $\mathcal{M}$. Instead of selecting the last or best frame for evaluation, POC uniformly samples multiple frames from $V$ at a frame rate of $r$ (denoted as $\text{POC}@r$), obtaining the frame set $\hat{V}_{r}=\{f_{k}\}_{k=1}^{N}$, where $N$ is the total number of sampled frames. To determine whether a frame adheres to the text condition, we employ a VLM as the judge model $\mathcal{J}$. Specifically, POC@r@r is formulated as:
@@
Table 2: Performance of mainstream video models arcoss 16 tasks within VIPER. We report the POC@@1.0 for each domain and summarize the average OC, POC and hacking rate. Best performance is bolded.
@@ 4.2 Hierarchical Rubric
To facilitate accurate consistency judgment, we design a hierarchical rubric for $\mathcal{J}$. Specifically, the rubric is structured into three distinct levels. Specific examples are provided in Appendix C.
@@
At the top tier, system prompt defines the task context, evaluation criteria, and methodology. Adopting the popular test-time scaling paradigm Guo et al. (2025a), the system prompt requires the model to first generate the analysis of process and outcome consistency within <think> tags, and subsequently provide the decision for PC, OC and POC within <answer> tags.
@@
In the middle, domain introduction provides a concise domain overview, highlighting the specific evaluation focus within the field. Crucially, all tasks belonging to the same domain share this unified introduction.
@@
At the instance level, task constraints imposes fine-grained evaluation criteria, which are reorganized from the process constraints $C$ and the reference $g$ collected previously.
Collectively, this hierarchical rubric covers from macro-level task definitions to micro-level requirements, providing comprehensive evaluation guidance. Moreover, this design is highly scalable. To support new tasks or domains, we only need to add the corresponding constraints and introductions.
@@ 5.2 Experiment Result
We present the main experiment results in Table 2, from which we summarize following observations. Firstly, all models struggle with VIPER, with POC@@1.0 scores consistently below 30%. Sora 2 performs the best, achieving a score of 23.3%, followed by Veo 3.1 and Wan 2.6, scoring 20.3% and 18.6%. The remaining models fail to surpass a score of 10%. This indicates that current video generation models are not yet capable of functioning as general visual reasoners.
Furthermore, severe outcome-hacking is observed across all models. Notably, Veo 3.1 experiences the most severe hacking, with a rate of 35.8%. This suggests that, while these models possess the necessary knowledge for fulfill reasoning goals, they struggle to faithfully represent the reasoning process through video frames. As such, improving process-level reasoning is a crucial step toward advancing model performance.
Lastly, the models demonstrate varying strengths across different tasks. Symbolic reasoning, in particular, proves to be a major challenge for most models, which struggle to generate accurate and legible text. However, Sora 2 excels in this domain, achieving a POC of 58.3%, which highlights its robust text rendering and reasoning capabilities. In contrast, for tasks like planning and physics which are more closely tied to real-world scenarios, Veo 3.1 outperforms Sora 2, underscoring its superior capabilities in real-world physics simulation.
Takeaway 1 Current video models are far from being universal visual reasoners, often reaching correct goal via invalid trajectories.
@@
Table 3: Effect of test-time scaling. We calculate Pass@@1, Pass@@ and Pass@@8 for Veo 3.1 and Sora2 with POC@@1.0.
@@ 5.3 Further Analysis
We investigate whether test-time scaling, effective in language reasoning, similarly benefits video reasoning. Pass@@$k$ is a widely used metric for evaluating language models, with recent studies indicating that performance can be significantly enhanced by increasing the number of samples $k$ Chen et al. (2021); Wang et al. (2023). To validate this in the GVR tasks, we sample multiple videos on VIPER using Sora 2 and Veo 3.1 and reporting Pass@@1, 4, and 8 with POC@@1.0 in Table 3. As illustrated, scaling the sampling times notably improves performance. For instance, the overall accuracy of Veo 3.1 increases from 20.3% to 43.2%, particularly in challenging domains such as symbolic reasoning. However, performance gains saturate as $k$ increases. Domains such as spatial and physics reasoning remain challenging, indicating that while test-time scaling provides a notable boost, it cannot fully compensate for reasoning limitations in current video models.
@@
Takeaway 2 While test-time scaling improves video model performance, it fails to resolve fundamental reasoning bottlenecks.
@@
To further validate the robustness of our evaluation paradigm, we compare the evaluation results from the judge model with those from human experts. As illustrated in Figure 4, by providing the judge model with our hierarchical rubric, its alignment with human evaluators generally exceeds 90%, whereas the raw evaluation achieves only about 60%. These results demonstrate the effectiveness of our rubric.
Takeaway 3 A hierarchical rubric is essential for reliable VLM-based evaluation, bridging the gap between automated metrics and human judgment where unstructured prompts fail.
Figure 4: Human-model alignment under combinations of raw input and hierarchical rubric.
Figure 5: Representative failure patterns observed from VIPER.
@@ 5.4 Failure Patterns
Upon observing generation results from VIPER, we identify several representative failure modes prevalent in current video models. We categorize these patterns below and illustrate specific cases in Figure 5, aiming to provide insights for the future optimization of video generation models.
@@
We observe that models often struggle to identify implicit constraints within tasks, such as not violating the maze structure when solving the maze or adhering to the movement rules in a checkmate-in-one scenario. To address this, we investigate whether explicitly prompting these constraints can mitigate the issue. We evaluate Veo 3.1 by supplementing the input with explicit process constraints $C$. As shown in Table 5, providing this auxiliary information significantly enhances model performance. This result indicates that the failure to respect constraints is not attributed to a deficiency in the model’s generation capabilities, but rather stems from an insufficient comprehension of the instructions.
Takeaway 4 The failures of video models are sometimes not due to a lack of fundamental generation abilities, but rather the difficulty in uncovering implicit intentions.
@@
Current video generation models typically produce outputs of fixed duration. However, reasoning tasks like checkmate-in-one often require short, precise sequences (e.g., 1-2 seconds). Ideally, the model should maintain a static frame upon task completion. Ideally, the model should maintain a still frame after completing the task. Instead, current models frequently continue to generate content after fulfill the task, introducing unnecessary motion or hallucinations that disrupt the integrity of the reasoning process.
@@
Unlike open-ended video generation, GVR tasks are highly structured and require precise, controlled editing. Models must strictly adhere to instructions without altering unrelated regions. However, we observe that current models struggle to define clear editing boundaries, often exhibiting “leakage” where background elements are altered, extra objects are introduced, or original task conditions are violated.
@@
The readability of text within the generated video is a crucial factor in determining reasoning quality, particularly in symbolic reasoning tasks where the model is expected to render the solution process in textual form. Currently, most models struggle to generate fluent and readable text passages, often producing garbled or incoherent glyphs. As a result, even if the reasoning outcome is correct, the video fails to provide a coherent and interpretable reasoning process.